OpenStudy (anonymous):

find the gradient of the straight line 3y=2x-11

OpenStudy (anonymous):

the gradient is the slope of a line so just write it in slope intercept form y=mx+b and m is ur slope since we have: 3y=2x-11 if we divide each side by 3 or formula will look like the slope intercept formula: y=(2/3)x-11/3 2/3=m & 11/3=b since we want m 2/3 is ur answer

hero (hero):

\[\color {green} {\huge\checkmark}\]